OpenStudy (anonymous):

3x^3y+12xy-3x^2y-12y rewrite this expression so the gcf is factored completely.

OpenStudy (anonymous):

\[x^3-x^2+4 x-4=(x-1) \left(x^2+4\right) \]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

my answer: gcf=3y

OpenStudy (anonymous):

\[3 x^3 y-3 x^2 y+12 x y-12 y=3 y (x-1) \left(x^2+4\right) \]

OpenStudy (anonymous):

that was part a can you help with part b and c?

OpenStudy (anonymous):

part b says to rewrite the expression completely factored and show steps part c wants to know would the outcome change if two of the middle terms of the problem are switched @robtobey